Circular No. 9062

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 222P/2009 MB_{9} = 2004 X1 (LINEAR) R. H. McNaught, Siding Spring Survey, writes that 2009 MB_{9} (cf.

MPEC 2009-M56; MPS 289318, 289477, 289905) appeared cometary on Aug. 2. H. Sato, Tokyo, suggested identity with comet P/2004 X1, the indicated correction to the prediction on MPC 56804 (erroneously cited as MPEC 56804 on CBET 1893) being ΔT = −2.2 days. The rediscovery observation is given below; further details are on MPEC 2009-P09. The comet has been assigned the permanent number 222P (MPC 66631).

2009 UT α_{2000} δ_{2000} Mag. Observer
June 29.47208 15^{h}03^{m}58.^{s}15 −38^{o}51′13′′.4 18.3 McNaught

COMET C/2009 P1 (GARRADD) R. H. McNaught, Siding Spring Observatory, reports the discovery by

G. J. Garradd (discovery observation below) of a tailless object with an approximately circular coma 15′′ across on 40-s CCD exposures with the 0.5-m Uppsala Schmidt telescope. Further observations were obtained by Garradd on the following night and at two other observatories in response to placement of the discovery on the NEO Confirmation Page.

2009 UT α_{2000} δ_{2000} Mag. Observer
Aug. 13.76794 0^{h}16^{m}01.^{s}27 −44^{o}33′23′′.1 17.5 Garradd

The available astrometry, the following very uncertain preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2009-P54.

T = 2011 Aug. 10.229 TT  ω = 95.^{o}562
Ω = 344.087 }2000.0
q = 1.25329 AU  i = 82.646

COMET P/2009 O3 (HILL) Additional astrometry of this comet (cf. IAUC 9058) shows it to be of

short period; orbital elements from MPEC 2009-Q05:

T = 2009 May 17.986 TT  ω = 154.^{o}293
e = 0.68496  Ω = 183.729 }2000.0
q = 2.44452 AU  i = 16.189
a = 7.75933 AU n^{o} = 0.045600 P = 21.61 years
2009 August 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green